# Optical excitation of bulk plasmons in n-doped InAsSb thin films : investigating the second viscosity in electron gas

### v1

@@ I.2 Resonances in a metallic slab
When the angle of incidence is different from zero, a vertical oscillation is predicted to occur at exactly the screened frequency $\omega_{0}$ when the electric field penetrates the entire slab. For a rigid gas to oscillate vertically, charges must accumulate on both surfaces of the slab, creating a restoring force that drives the oscillation and induces a resonance. This mechanism has been suggested by Ferrell initially[3]. Shortly thereafter, a resonance has been observed more or less in the predicted conditions[18] and it is sometimes said that it corresponds to the excitation of a so called Brewster mode. There is however no guided or cavity mode that is excited in that case. The nature of the resonance is fundamentally different from a cavity resonance for light, or of the excitation of a guided mode.
When interactions between electrons are considered, a different picture emerges[5, 6], which can be considered more accurate and more complete. Surface charges can no longer be described as such; instead, they appear as volume charges near the surface, a phenomenon known as ’smearing’ of the charges. The electron gas can no longer be viewed as a rigid entity moving as a whole; rather, it supports plasmons.
When a plasmon is excited by incident light, the wavevector component along the interface is conserved. However, the plasmon’s wavevector is much larger than that of light, as discussed earlier. As a result, the plasmon’s wavevector is dominated by its vertical component, regardless of the incidence angle. The slab can therefore be treated as a cavity for these waves, with resonances occurring at relatively small thicknesses. Melnyk and Harrison’s calculations[5] demonstrate that only odd-order resonances can be excited. Although the fundamental resonance is more pronounced and was initially thought to confirm Ferrell’s mechanism, experiments —first with potassium in the 1970s[19], and more recently with thin films of highly doped semiconductors[8] — have observed higher-order resonances that are not consistent with Ferrell’s interpretation. This confirms that the metallic slab behaves as a plasmon cavity, supporting odd resonances, and that a spatially dispersive model is required for an accurate description of its response.
@@ II.1 Sample preparation
We have prepared two series of samples of Si doped n-InAsSb, which we will denote #1 and #2. The two series differ by their doping level. Each series derives from a single 204 nm thick n-InAsSb which is subsequently etched to guarantee both an homogeneous level of doping in each series and a good control of the thickness.
@@
The n-InAsSb layer is a digital alloy, i.e. a short period superlattice, which allows the deposited thickness, 204 nm, to be accurately monitored for both samples. The doping level of the Si doped n-InAsSb layer, estimated using the optical method described in [4], is of $5.3\times 10^{18}$ cm-3 and $1.20\times 10^{19}$ cm-3 for sample #1 and #2, respectively. The determination of the doping level, as mentioned above, is a reflection experiment under polarized light with an angle of incidence of 60°. A home-made abacus then allows to deduce the carrier density from the measurement of the resonance wavelength.
The initial samples are broken into several pieces and each one is then etched to obtain a series of samples, whith a thickness varying from 7 nm to 204 nm. The samples are wet etched at 20°C in a 2:1 ratio of citric acid ($C_{6}H_{8}O_{7}$) and hydrogen peroxide ($H_{2}O_{2}$) solution, the time of exposition allowing to control the final thickness. To evaluate the final thickness of each sample, a part of the surface is protected from wet etching by a small drop of photoresist (AZMIR-701) deposited on one edge and heated at 90°C for 1 min 30 s. After removal of the photoresist, the remaining step is measured by atomic force microscopy (AFM) and subtracted from the total thickness of 204 nm. The resulting thicknesses, despite similar exposure time, differ for series #1 and #2.
@@ II.2 Optical response measurement
Figure 2: Reflectance spectrum under p-polarized light with an angle of incidence of 60° (black line) and pseudo-ATR spectrum of sample #1 with a thickess of 204 nm.
This resonance can be excited either in reflection, when illuminated from air, or when illuminated from inside a germanium crystal using a Schwarzschild objective, mounted on a microscope (Hyperion 3000) and coupled to an FTIR spectrometer (Bruker, Vertex 70). Given the high index of the Ge crystal (around 4), such a configuration amounts to exciting the structure using a prism, with an angle of incidence between 21.5° and 37°, determined by the objective (see Fig.3). For media with a refractive index smaller than 1.6, this means total internal reflection will occur and this would constitute a attenuated total internal reflection (ATR) setup. However, especially in the mid-IR range, materials most often present larger indexes so that we call this configuration ”pseudo-ATR” in the rest of the paper.
@@
Figure 3: Experimental setup in the pseudo-ATR configuration. The sample is illuminated with incidence angle ranging from 21 to 37.
@@ II.3 Plasmon resonances
Using the samples of different thicknesses, we first map the resonances that can be seen on the different spectra shown in Figure 4 and Figure 5 and compare the position with predictions provided using a hydrodynamic model.
The pseudo-ATR spectra of series #1 are shown in Figure 4 in log scale. Spectra are background corrected with a reference sample of GaSb with a 210 nm thick buffer layer of undoped GaSb. In each spectrum the fundamental plasmon resonance (that we often call Ferrell resonance) can be seen at short wavelength and it progressively blueshifts when the thickness decreases. We identify this peak as the first order plasmon mode, $\ell=1$. At higher wavelengths, some additional peaks corresponding to the odd high order plasmon modes, $\ell=3,5,7,...$ can be observed, even if higher order resonances can be difficult to identify. All these resonances also blueshift with decreasing thickness. Such a behaviour can be expected for plasmon resonances as will be explained below. Such a blueshift is however not found for the Ferrell resonance in the framework of the Drude model. Note that the small peak at 1250 cm-1 is due to the GaSb substrate. To accurately extract the wavenumber of each high order bulk plasmon mode, we first fit the first order bulk plasmon with a Lorentzian function, which we subtract to the experimental data.
Figure 4: Pseudo-ATR spectra for several thickness of sample 1#.
The position of the resonances as a function of the sample thickness is summarized in Figure 5. The colored disks correspond to the wavelength of the bulk plasmon modes, white, red, blue and pink for $\ell=1,3,5$ and $7$ respectively. Error bars on the disk position are indicated with vertical and horizontal lines.
Figure 5: Plot of the simulated spectra of log10((RHDM(s) + RHDM(p))/2) where RHDM(s) and RHDM(p) are the pseudo-ATR spectra under s- and p-polarizations respectively. The colored disks correspond to position of the plasmon resonances as determined using the experimental data. We identify each resonance with an odd number. Error bars are the vertical and horizontal ticks for each disk.
All bulk plasmon modes blueshift with decreasing thickness. This result is confirmed by the simulation of the spectra using the HDM described above. The color map in Figure 5 corresponds to log10((RHDM(s) + RHDM(p))/2) where RHDM(s) and RHDM(p) are respectively the pseudo-ATR spectra under s- and p-polarizations. Experiments and simulations agree relatively well. We have used the same parameters for both samples to model the n-type GaSb substrate and the GaSb buffer. The n-type GaSb substrate is correctly modelled with a Drude model using parameters $\varepsilon_{\infty}=14.4$, $\omega_{p}=2.2\times 10^{14}$ rad.s-1, and $\gamma=5\times 10^{12}$ rad.s-1. For the GaSb buffer we used the refractive index in[20]. For the n-InAsSb layer, the fitting parameters of the HDM are the carrier density, $n_{0}=5.5\times 10^{18}$ cm-3, the damping parameter, $\gamma=6.4\times 10^{12}$ rad.s-1 and the hydrodynamic parameter, $\beta=8.95\times 10^{5}$ m.s-1.
@@
The same study is carried out on sample #2. Figure 6 is the pseudo-ATR spectra for different thicknesses. As the sample presents a higher level of doping, the first order of the bulk plasmon mode is blueshifted compared to sample #1. The resonances are also broadened. However, it is possible to observe the high order bulk plasmon modes up to $\ell=11$. Once all the odd orders of the bulk plasmon mode were extracted, we carried out a fit with the HDM to get the parameters of the sample #2 (Fig. 7).
Figure 7: Plot of the simulated spectra of log10((RHDM(s) + RHDM(p))/2). The colored disks correspond to the order of the bulk plasmon modes. Error bars are the vertical and horizontal ticks for each disk.
The behaviour of the bulk plasmon modes is the same as for sample #1. Once again, the experiments and simulations agree quite well. We have used the following fitting parameters of the HDM to describe the n-InAsSb layer: the carrier density, $n_{0}=1.2\times 10^{19}$ cm-3, the damping parameter, $\gamma=8\times 10^{12}$ rad.s-1 and the hydrodynamic parameter, $\beta=9.29\times 10^{5}$ m.s-1.
@@ III Improved parameter retrieval
As demonstrated by the results above, the pseudo-ATR configuration significantly enhances the response of semiconductor slabs, allowing us to accurately determine the parameters of the hydrodynamic model, which is usually challenging[21]. To streamline this process, we have developed an automated method to retrieve these parameters. Our method includes three key components. Firstly, we use the hydrodynamic model with a complex hydrodynamic parameter $\beta$, which will be explained in detail below. Secondly, we define a cost function that prioritizes matching the overall shape of the curves rather than just minimizing the distance between data points, aligning better with the intuitive assessment of physicists. Thirdly, we employ the Differential Evolution (DE) global algorithm, known for its efficiency in solving physical problems[22, 23], ensuring that we do not miss any satisfactory parameter values.
@@
We first chose to fit the experimental data using parameters that rely minimally on underlying assumptions and can be considered empirical, such as $\chi_{b}$, $\omega_{p}$, and $\gamma$ for the Drude model, and a complex value of $\beta$ for the hydrodynamic part. To accurately fit the experimental data, we also include a parameter for the background signal and another for scaling the data.
In addition, we have fitted our data with a dispersive imaginary part for the complex value of $\beta^{2}$, which will be justified thoroughly below. The complex value of $\beta^{2}$ is denoted by $\eta^{2}$ and can be expressed as:
@@
Here $\xi$ is an additional parameter whose origin will be explained later.
@@
where $\lambda_{i}$ is a wavelength for which a reflectance $R(\lambda_{i})$ has been measured, $\Pi$ represents the parameters of the simulation ($\chi_{b}$, $\omega_{p}$, $\gamma$,$\beta$,$\frac{\xi}{n_{0}\,m^{*}}$,background signal intensity, scaling parameter), and $S(\Pi,\lambda_{i})$ is the reflectance of the structure simulated for the given parameters and wavelength.
@@
In order to find the minimum of the cost function, we use Differential Evolution (DE) (more precisely its Quasi-Oppositional version[23]) to look for satisfactory values of the parameters. DE is a non-deterministic algorithm and can be expected to provide different results each time it is run. We also take the results of DE and then perform a steepest descent[23], in order to be sure we have reached a local minimum.
The numerical computation of the reflectance, taking into account an imaginary part for $\beta^{2}$, has been included in the open source PyMoosh software[24]. In addition, we have made examples available with the code under the form of a Jupyter notebook, leveraging the features of PyMoosh to retrieve the parameters of the model efficiently[25].
The resulting fits are particularly satisfactory despite the uncertainty in the incidence angle and the lack of any averaging. An example is shown Fig. 9 a 70 nm thick sample from series #2, where the plasmon resonances are clearly visible. These fits have been obtained by assuming a constant imaginary part for $\eta^{2}$. This shows how accurately the hydrodynamic model with a complex value for $\eta^{2}$ is able to model the response of the semiconductor.
Figure 9: Reflectance as a function of wavelength for a semiconductor slab from series #2 with a thickness of 70 nm. The observed plasmon resonances align closely with the model, demonstrating the accuracy of the hydrodynamic fit using a complex $\eta^{2}$ with an imaginary part (the imaginary part is constant here).
@@
Figure 10: Resonance characterized by $\ell=3$ for a 106 nm thick sample of series #1, for different values of $\frac{\xi}{n_{0}m^{*}}$ ranging from 0 to 7×$\times 10^{-4}$ in m2.s-1. The orange curve corresponds to $\xi=0$. When the value of $\xi$ increases, the resonance profile becomes smoother and deviates from a Fano profile.
@@
The parameter retrieval can be performed on a single spectrum for a given thickness or across multiple thicknesses. Based on our experience, a fit on a single spectrum is sufficient when the main resonance and at least one higher-order resonance are clearly visible. The more resonances that are observable, the better the determination of $\beta$. For sample #2, selecting a spectrum with enough resonances yielded the most reliable results. For instance, using the spectrum for a 100 nm thickness (shown in Fig. 11), we obtained estimations of $n_{0}=1.39\times 10^{19}$ cm-3, $m^{*}=0.077\,m_{0}$ and $\xi=5.7\times 10^{-10}$ Pa.s.
@@
Figure 11: Experimental reflectance data for a 110 nm thick InAsSb layer of sample #2 in the pseudo-ATR configuration, compared with the best fit obtained for the given model parameters. The optimized fit is displayed only within the wavelength range used for optimization.
@@
However, to get more reliable values, a fit taking into account all the thicknesses can be preferable. At small thicknesses, the shift of the main resonance away from $\omega_{0}$ is directly linked to the value of $\beta$, thus providing important information for the optimization. For sample #1, the results are very convincing, as shown Fig. 12. The values obtained are $n_{0}=5.31\times 10^{18}$ cm3, $m^{*}=0.057\,m_{0}$ and $\xi=2.11\times 10^{-10}$ Pa.s.
We underline that the optimization terminates with a steepest descent but still provides slightly different results each time. This means the cost function has many local minima and that the problem, from the optimization point of view, is challenging. This produces uncertainty in the values provided for the parameters of the model, and consequently for the estimation of $n_{0}$, $m^{*}$ and $\xi$.
Especially, even small variations in the estimation of $\beta$ result in a relatively large (around 10%) variation in the estimation of $n_{0}$. Conversely, the value provided for the effective mass of the electron is much more stable. Comparing these results with a manual retrieval as shown Sec. II.3 is a good idea, but we are confident given the present results that with a more controlled incidence angle or a more precise determination of the thickness of each sample, the accuracy of the whole process can be improved.

### v2

@@ I.4 Resonances in a metallic slab
Within the framework of the Drude model, when a horizontal slab containing a free electron gas is illuminated in normal incidence, the gas is free to oscillate horizontally. No resonance is associated with such an oscillation. When the angle of incidence is different from zero, a vertical oscillation is predicted to occur at exactly the screened frequency $\omega_{0}$ when the electric field penetrates the entire slab. For a rigid gas to oscillate vertically, charges must accumulate on both surfaces of the slab, creating a restoring force that drives the oscillation and induces a resonance – a mechanism suggested by Ferrell[ferrell1962plasma]. Shortly thereafter, a resonance was observed more or less in the predicted conditions[steinmann1960experimental] and it is sometimes said that it corresponds to the excitation of a so called Brewster mode. There is however no guided or cavity mode that is excited in that case. The nature of the resonance is fundamentally different from a cavity resonance for light, or of the excitation of a guided mode.
When interactions between electrons are considered, surface charges can no longer be described as such; instead, they appear as volume charges near the surface, a phenomenon known as ’smearing’ of the charges. The electron gas can no longer be viewed as a rigid entity moving as a whole; rather, it supports plasmons. When a plasmon is excited by incident light, the wavevector component along the interface is conserved. However, the wavevector of the plasmon being much larger than that of light, it is dominated by its vertical component, regardless of the incidence angle. The slab can therefore be treated as a cavity for these waves, with resonances occurring at relatively small thicknesses as shown by Melnyk and Harrison[melnyk1968resonant, melnyk1970theory].
With that picture in mind, it is easy to estimate the position of the resonances. Whatever the incidence angle, the plasmon can be considered to propagate almost perfectly perpendicularly to the interfaces so that the resonance condition for such a cavity can be expressed as:
$k=\frac{\ell\pi}{h}$ (14)
where $\ell$ is a non-zero integer, $h$ the thickness of the slab and $k$ the plasmon wavevector.
Using the lossless dispersion relation for the plasmon, this provides an approximation for the resonance frequencies:
$w_{m}=\sqrt{w_{0}^{2}+\left(\frac{\ell\beta\pi}{h}\right)^{2}}.$ (15)
Despite its simplicity and the fact that the losses are neglected, this formula is quite accurate, as shown in Fig. 2, where the lossless dispersion relation is used to compute the position of the resonance on the spectrum. The spectrum shown is computed with $\xi=0$ to better see the position of the resonances. The figure shows clearly that only the resonances corresponding to an odd value of $\ell$ can be seen, as already noticed by Melnyk and Harrisson[melnyk1968resonant] who did not provide any explanation, though.
The cavity picture allows us to understand why the first resonance (the Ferrel resonance) can in fact not occur at precisely $\omega_{0}$. As shown in Figure 2, there is a shift between the position of the first resonance predicted by the hydrodynamic model and by a Drude model, which assumes a rigid electron gas. While this shift is small for such a thick structure, it becomes much larger for thinner structures. Such a large shift is thus a clear sign of a nonlocal response, dictated by the dispersion relation of plasmons. It gives a clear indication on the value of $\beta$. This simple model also allows us to understand that, when $\beta$ tends to zero, all the resonances coalesce into a single resonance, corresponding to the Ferrel resonance within the framework of the Drude model.
Figure 2: Top: Simulated reflectance spectrum for a 70 nm thick slab without any imaginary part for $\beta^{2}$ (blue curve) and prediction using the Drude model only (orange curve). Only the odd resonances (grey dashed lines) are visible, while even resonances are absent. A simple cavity formula provides an accurate way to compute the position of these resonances. Bottom: Dispersion curve for plasmons, and cavity resonance conditions $k=\frac{m\pi}{h}$ allowing to compute the resonance frequencies shown at the top of the figure.
The fact that the even resonances of a cavity cannot be excited is a phenomenon that typically arises when a cavity is excited in phase and with the same amplitude from both sides, a phenomenon sometimes referred to as coherent perfect absorption [baranov2017coherent] or more accurately as the interferometric enhancement of the absorption [lemaitre2017interferometric].
At first glance, the physical situation here does not seem to fit this description because the structure is illuminated from above only. However, the slab functions as a two-mode cavity[pitelet2017fresnel]. The transverse wave also traverses the semiconductor slab, and since we are very close to $\omega_{0}$, the effective refractive index approaches zero, resulting in a vanishing wavevector and a diverging wavelength. To be precise, at the considered incidence angle, the transverse wave is usually evanescent in the medium but has a very large penetration depth. In both scenarios, the slab thickness is small compared to either the wavelength or the penetration depth, allowing it to reach the bottom interface while maintaining the same phase and amplitude. The plasmon that is generated at the bottom interface is then expected to have the same phase as the plasmon generated at the top interface and a comparable amplitude, thus canceling the even resonances of the cavity.
We underline that this also leads to the enhancement of odd resonances. This is fundamental in the so-called coherent perfect absorption and explains the large absorption cross-section of plasmonic nanocavities[moreau2012controlled] because it enhances absorption at resonance by a factor of typically four. Without such a phenomenon, it is likely only the main resonance could have been observed here.
Figure 3: Field profiles confirming selective excitation of odd plasmon modes. Top: Modulus of $E_{z}$ and real part of the charge density for th fundamental mode. Bottom: Same quantities for the mode characterized by $\ell=3$. The antisymmetric charge distribution for the fundamental mode explains the restoring force responsible for the main resonance.
To verify the cavity picture, we extended the PyMoosh simulation library[langevin2024pymoosh] to handle nonlocal response using a generalized S-matrix formalism[benedicto2015numerical] and compute charge density distributions within the semiconductor. Figure 3 displays the spatial profiles at resonance frequencies, showing the characteristic odd symmetry of the excited modes. The charge density extends deep into the bulk rather than being confined to interfaces as in the Drude model, confirming the bulk plasmon nature of these resonances. This visualization directly demonstrates why even modes remain dark: the symmetric excitation from both interfaces naturally selects odd-symmetry modes.
@@ II.1 Sample preparation
We prepared two series of samples of Si doped n-InAsSb, which we will denote #1 and #2. The two series differ by their doping level. Each series derives from a single 204 nm thick n-InAsSb which is subsequently etched to guarantee both an homogeneous level of doping in each series and a good control of the thickness.
@@
The n-InAsSb layer is a digital alloy, i.e. a short period superlattice, which allows the deposited thickness, 204 nm, to be accurately monitored for both samples. The doping level of the Si doped n-InAsSb layer, estimated using the optical method described in [taliercio2014brewster], is of $5.3\times 10^{18}$ cm-3 and $1.20\times 10^{19}$ cm-3 for sample #1 and #2, respectively. While surface accumulation layers are well-documented in InAs-based materials [wieder1974transport, casias2019carrier, rogalski2020inassb], with typical densities of $\sim 10^{12}$ cm-2 (corresponding to $\sim 10^{18}$ cm-3 over a few nm), our intentional doping levels are significantly higher. At these high bulk doping concentrations, any surface effect would manifest as charge depletion rather than accumulation, which would shift the plasma frequency to longer wavelengths—opposite to the spectral features observed experimentally. Therefore, we can reliably consider the doped layer as homogeneous for our optical analysis. The determination of the doping level, as mentioned above, is a reflection experiment under polarized light with an angle of incidence of 60°. A home-made abacus then allows to deduce the carrier density from the measurement of the resonance wavelength[taliercio2014brewster].
The initial samples are broken into several pieces and each one is then etched to obtain a series of samples, with a thickness varying from 7 nm to 204 nm. The samples are wet etched at 20°C in a 2:1 ratio of citric acid ($C_{6}H_{8}O_{7}$) and hydrogen peroxide ($H_{2}O_{2}$) solution, the time of exposition allowing to control the final thickness. To evaluate the final thickness of each sample, a part of the surface is protected from wet etching by a small drop of photoresist (AZMIR-701) deposited on one edge and heated at 90°C for 1 min 30 s. After removal of the photoresist, the remaining step is measured by atomic force microscopy (AFM) and subtracted from the total thickness of 204 nm. The resulting thicknesses, despite similar exposure time, differ for series #1 and #2.
@@ II.2 Optical response measurement
Figure 4: Reflectance spectrum under p-polarized light with an angle of incidence of 60° (black line) and pseudo-ATR spectrum of sample #1 with a thickness of 204 nm.
This resonance can be excited either in reflection, when illuminated from air, or when illuminated from inside a germanium crystal using a Schwarzschild objective, mounted on a microscope (Hyperion 3000) and coupled to an FTIR spectrometer (Bruker, Vertex 70). Given the high index of the Ge crystal (around 4), such a configuration amounts to exciting the structure using a prism, with an angle of incidence between 21.5° and 37°, determined by the objective (see Fig.5). For media with a refractive index smaller than 1.6, this means total internal reflection will occur and this would constitute an attenuated total internal reflection (ATR) setup. However, especially in the mid-IR range, materials most often present larger indexes so that we call this configuration ”pseudo-ATR” in the rest of the paper.
@@
Figure 5: Experimental setup in the pseudo-ATR configuration. The sample is illuminated with incidence angle ranging from 21 to 37. The Ge prism is in direct contact with the sample, but given the incidence angles, no total internal reflection occurs.
@@ II.3 Plasmon resonances
Using samples of different thicknesses, we systematically map the plasmon resonances visible in our pseudo-ATR configuration. The enhanced visibility afforded by this setup allows us to observe resonances up to order $\ell=11$ for sample #2, far exceeding previous reports[de2018viscoelastic]. While these high-order resonances appear at longer wavelengths due to our thicker samples and lower doping levels, they demonstrate that bulk plasmon physics is not confined to ultra-thin films but persists in semiconductors nearly an order of magnitude thicker.
The pseudo-ATR spectra of series #1 are shown in Figure 6 in log scale. Spectra are background corrected with a reference sample of GaSb with a 210 nm thick buffer layer of undoped GaSb. In each spectrum the Ferrel resonance can be seen at short wavelength and it progressively blueshifts when the thickness decreases. This blueshift cannot be explained within the Drude model where the Ferrell resonance position is thickness-independent, providing clear evidence of nonlocal effects described by the plasmon dispersion relation.
We identify this peak as the first order plasmon mode, $\ell=1$. At higher wavelengths, some additional peaks corresponding to the odd high order plasmon modes, $\ell=3,5,7,$... can be observed, even if higher order resonances can be difficult to identify. All these resonances also blueshift with decreasing thickness. Note that the small peak at 1250 cm-1 is due to the GaSb substrate. To accurately extract the wavenumber of each high-order plasmon resonance, we first fit the first-order plasmon resonance with a Lorentzian function, which we subtract from the experimental data. This enhances the visibility of the high order resonances.
Figure 6: Pseudo-ATR reflectance spectra for several thickness of sample #1.
The position of the resonances as a function of the sample thickness is summarized in Figure 7. The colored disks correspond to the wavelength of the plasmon resonances, white, red, blue and pink for $\ell=1,3,5$ and $7$ respectively. Error bars on the disk position are indicated with vertical and horizontal lines.
Figure 7: Position of the plasmon resonances as determined using the experimental data for sample #1. We identify each resonance with an odd number. Error bars are the vertical and horizontal ticks for each disk.
All plasmon resonances clearly blueshift with decreasing thickness, following the cavity resonance condition $k=\ell\pi/h$ combined with the plasmon dispersion relation, as predicted by Eq. 15.
@@
The same study is carried out on sample #2. Figure 8 is the pseudo-ATR spectra for different thicknesses. As the sample presents a higher level of doping, the first order of the bulk plasmon mode is blue-shifted compared to sample #1. The resonances are also broadened. However, it is possible to observe the high order bulk plasmon modes up to $\ell=11$. To our knowledge, this represents the highest-order bulk plasmon resonance reported in doped semiconductors, made possible by the symmetric refractive index configuration and the relatively low losses in our samples.
Figure 9: Position of the plasmon resonances as determined using the experimental data for sample #2. We identify each resonance with an odd number. Error bars are the vertical and horizontal ticks for each disk.
The data provide two key signatures for parameter extraction: (i) the blueshift of resonances away from $\omega_{0}$, which directly determines Re($\beta$) via Eq. 15, and (ii) the absence of the sharp Fano profiles predicted for real $\beta$ (cf. Fig. 2), requiring an imaginary part to explain the observed smoothing. These features are sufficiently distinctive that even a single thin sample with visible high-order resonances contains enough information to extract all the model parameters as demonstrated in the following section.
@@ III Improved parameter retrieval
To streamline the parameter retrieval process, we have developed an automated method to retrieve these parameters. Our method includes three key components. Firstly, we use the hydrodynamic model with a complex hydrodynamic parameter $\beta$, which will be explained in detail below. Secondly, we define a cost function that prioritizes matching the overall shape of the curves rather than just minimizing the distance between data points, aligning better with the intuitive assessment of physicists. Thirdly, we employ the Differential Evolution (DE) global algorithm, known for its efficiency in solving physical problems[barry2020evolutionary, bennet2024illustrated], ensuring that we do not miss any satisfactory parameter values.
@@
We first chose to fit the experimental data using parameters that rely minimally on underlying assumptions and can be considered empirical, such as $\chi_{b}$, $\omega_{p}$, and $\gamma$ for the Drude model, and a complex value of $\eta^{2}$ for the hydrodynamic part:
@@
We actually introduce $\tau=\frac{\xi}{n_{0}m^{*}}$ as a parameter to be optimized and retrieve $\xi$ afterwards. Finally, we also include a parameter for the background signal and another for scaling the data.
@@
where $\lambda_{i}$ is a wavelength for which a reflectance $R(\lambda_{i})$ has been measured, $\Pi$ represents the parameters of the simulation ($\chi_{b}$, $\omega_{p}$, $\gamma$, $\beta$, $\tau=\frac{\xi}{n_{0}\,m^{*}}$, background signal intensity, scaling parameter), and $S(\Pi,\lambda_{i})$ is the reflectance of the structure simulated for the given parameters and wavelength.
@@
In order to find the minimum of the cost function, we use Differential Evolution (DE) (more precisely its Quasi-Oppositional version and then a steepest descent to refine the results[bennet2024illustrated]) to look for satisfactory values of the parameters.
The numerical computation of the reflectance, taking into account an imaginary part for $\beta^{2}$, has been included in the open source PyMoosh software[langevin2024pymoosh]. In addition, we have made the code we used available under the form of a Jupyter notebook, we is made possible because PyMoosh combines simulation and optimization tools[latest].
@@
[…]
@@
The parameter retrieval can be performed on a single spectrum for a given thickness or across multiple thicknesses. Based on our experience, a fit on a single spectrum is sufficient when the main resonance and at least one higher-order resonance are clearly visible. The more resonances that are observable, the better the determination of $\beta$. For sample #2, selecting a spectrum with enough resonances yielded the most reliable results. For instance, using the spectrum for a 100 nm thickness (shown in Fig. 12), we obtained estimations of $n_{0}=1.40\pm 0.08\times 10^{19}$ cm-3, $m^{*}=0.083\pm 0.006\,m_{0}$ and $\xi=3.7\pm 1.0\times 10^{-10}$ Pa.s.
@@
Figure 12: Experimental reflectance data for a 70 nm thick InAsSb layer of sample #2 in the pseudo-ATR configuration, compared with the best fit obtained for the given model parameters. The predictions of the Drude model with the same parameters are also shown, to illustrate the necessity of the nonlocal description.
@@
However, to get more reliable values, a fit taking into account all the thicknesses can be preferable. At small thicknesses, the shift of the main resonance away from $\omega_{0}$ is directly linked to the value of $\beta$, thus providing important information for the optimization. For sample #1, the results are very convincing, as shown Fig. 13. The values obtained are $n_{0}=5.31\pm 0.5\times 10^{18}$ cm3, $m^{*}=0.057\pm 0.05\,m_{0}$ and $\xi=2.11\pm 0.6\times 10^{-10}$ Pa.s.
Uncertainty quantification for this inverse problem presents unique challenges that standard methods do not fully address. While multiple optimization runs consistently converge to similar parameter values (with a  5% variability typically), suggesting robust convergence despite the presence of multiple local minima, this statistical variability underestimates the true parameter uncertainty.
We therefore adopted a physics-informed sensitivity analysis: parameters were varied from their optimal values until the fit quality degraded beyond what would be considered acceptable. This approach yields conservative uncertainty estimates of approximately 10% for $n_{0}$ and $m^{*}$, capturing not only numerical variability but also the inherent ambiguity in defining acceptable fit quality given experimental noise and systematic effects. The uncertainty on the bulk viscosity $\xi$ is larger, as the resonance profile shows a gradual dependence on this parameter (see Fig. 11), making precise determination more challenging.
While rigorous uncertainty estimation for such inverse problems remains an open challenge, we anticipate that improvements to the experimental setup—particularly better control of the incidence angle distribution and more accurate thickness measurements—would significantly reduce these uncertainties before more sophisticated analysis methods become necessary. Indeed, the current approach of fitting multi-angle experimental data with plane-wave simulations already introduces systematic approximations that likely dominate over numerical uncertainties.